OpenStudy (anonymous):

a train travels 180 miles in the same time that a car travels 120 miles, the speed of the train is 20 miles per hour faster than the speed of the car find the speed of the train and the speed of the car?

OpenStudy (e.cociuba):

Train DATA: distance = 180 miles ; rate = x+20 mph ; time = d/r = 180/(x+20) hrs Car DATA: distance = 120 miles ; rate = x mph ; time = d/r = 120/x hrs time = time 180/(x+20) = 120/x Divide both sides by 60 to get: 3/(x+20) = 2/x 3x = 2(x+20) x = 40 mph (speed of the car) x+20 = 60 mph (speed of the train)